OpenStudy (anonymous):

Is this arithmetic or geometric ? \( {4n^2} \) Not sure how to approach this one. Do I \(4(n-1)^2 \)

OpenStudy (anonymous):

This has to do with Sequences

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

it's neither

OpenStudy (anonymous):

How do I check? Do I just 4*1^2 =16 4*2^2 = 64 and then check the for the difference and ratio ?

OpenStudy (anonymous):

you can do 3 consecutive terms to check. \[a_1=4(1)^2=4\] \[a_2=4(2)^2=16\] \[a_3=4(3)^2=36\] If \(\frac{ a_2 }{ a_1 }=\frac{ a_3 }{ a_2 }\) then it's geometric. If \(a_3-a_2=a_2-a_1\), then it's arithmetic

OpenStudy (anonymous):

Neither of those is true so the sequence is neither arithmetic or geometric

OpenStudy (anonymous):

Ok, that is what I was thinking. Thank you and I messed up on my on my powers. I was thinking (4*1)^2 for some reason but it is the way you have it.

OpenStudy (anonymous):

no problem. the geometric sequences will look like exponential functions and arithmetic sequences look like linear functions. If it's a higher order polynomial or some other type of function it's neither.